OpenStudy (anonymous):

Find the indicated probability using the standard normal distribution. P(z< -0.33 or z> 0.33)

OpenStudy (mathmate):

Hint: P(z< -0.33 or z> 0.33) =1-P(-0.33<z<0.33 =1-2P(0<z<0.33)